### v1

# ViTamin: Designing Scalable Vision Models in the Vision-Language Era

Abstract
Recent breakthroughs in vision-language models (VLMs) start a new page in the vision community. The VLMs provide stronger and more generalizable feature embeddings compared to those from ImageNet-pretrained models, thanks to the training on the large-scale Internet image-text pairs. However, despite the amazing achievement from the VLMs, vanilla Vision Transformers (ViTs) remain the default choice for the image encoder. Although pure transformer proves its effectiveness in the text encoding area, it remains questionable whether it is also the case for image encoding, especially considering that various types of networks are proposed on the ImageNet benchmark, which, unfortunately, are rarely studied in VLMs. Due to small data/model scale, the original conclusions of model design on ImageNet can be limited and biased. In this paper, we aim at building an evaluation protocol of vision models in the vision-language era under the contrastive language-image pretraining (CLIP) framework. We provide a comprehensive way to benchmark different vision models, covering their zero-shot performance and scalability in both model and training data sizes. To this end, we introduce ViTamin, a new vision models tailored for VLMs. ViTamin-L significantly outperforms ViT-L by 2.0% ImageNet zero-shot accuracy, when using the same publicly available DataComp-1B dataset and the same OpenCLIP training scheme. ViTamin-L presents promising results on 60 diverse benchmarks, including classification, retrieval, open-vocabulary detection and segmentation, and large multi-modal models. When further scaling up the model size, our ViTamin-XL with only 436M parameters attains 82.9% ImageNet zero-shot accuracy, surpassing 82.0% achieved by EVA-E that has ten times more parameters (4.4B).
1 Introduction
Figure 1: Practices of designing scalable vision models in the vision-language era. We benchmark modern vision models with various model and data scales under CLIP setting using DataComp-1B [30], leading to findings about data and model scalability, feature resolution, and hybrid architecture, which motivate us to develop ViTamin for VLM. ViTamin-L achieves superior zero-shot performance over ViT-L/14 [60] on ImageNet [86] and average 38 datasets [30], and advances a suite of 22 downstream tasks for Open-Vocabulary (OV) detection [111] and segmentation [124], and Large Multi-modal Model (LMM) tasks [67].
The past decades have witnessed significant progress in computer vision, like visual recognition tasks. The advent of AlexNet [53] marked a significant milestone, catalyzing the extensive evolution and dominance of Convolutional Neural Networks (ConvNets) [55, 37, 46, 73, 8, 32, 38, 72] in computer vision. More recently, with the debut of Vision Transformer [104, 23], a growing number of transformer-based architectures [108, 121, 71, 18, 116, 103] have shown great potential to surpass the prior ConvNet counterparts.
The rapid advancement of neural network design in computer vision can be attributed to a combination of factors. Among them, an important factor is the well-established benchmarks, allowing the community to examine the developments in a standardized way. Particularly, ImageNet [86] has become the de facto testing ground for new vision models. It not only sets a standard benchmark for visual recognition, but also serves as a mature pre-training dataset for transferring the network backbone to a variety of downstream tasks (e.g., detection and segmentation) [66, 73, 9, 38, 10, 51, 15, 107, 122, 123, 95, 117].
Recently, the emergence of vision-language models (VLMs) [82, 50] has changed the paradigm by leveraging the pre-training schedule on the extremely large scale noisy Internet data up to billions of image-text pairs [89], much larger than the ImageNet scale. VLMs not only produce strong and generalizable features [82, 50], but also excel in zero-shot downstream tasks [85, 77, 31, 54, 124, 68, 137]. However, unlike the ImageNet benchmark, where many types of neural networks are designed and blossomed [53, 91, 37, 46, 44, 97], the existing VLMs mostly employ the vanilla Vision Transformer (ViT) architecture [23] ***with only a few exceptions, e.g., ConvNeXt [72] by OpenCLIP [48]., and the recent benchmark DataComp [30] focuses on the data curation under the common (yet unverified) belief that ViTs scale much better than any other architectures in this vision-language era [20, 61] and thus ViT is all we need.
The current trend can be characterized by several key observations: (1) The high computational demand, requiring extensive resources [48] for months, is a significant barrier for advancing VLMs [82], limiting exploring diverse vision models. (2) Traditional vision models are mainly optimized for the ImageNet benchmark, which may not scale well for larger datasets [89, 30], unlike purely transformer-based architectures [104] that have proven scalable in language tasks [81, 102] and are now being adopted for VLMs as image encoders [23]. (3) Current VLM benchmarks focus on zero-shot classification/retrieval tasks [30], with a notable lack of downstream tasks involving open-vocabulary dense prediction [129, 31, 115, 135, 54, 22, 114, 124, 21], as well as a gap in assessing Large Multi-modal Models (LMMs) [57, 68, 67, 137].
In this paper, we aim to address the aforementioned issues with practices as shown in Fig. 1. To begin with, we establish a new test bed for designing vision models under the CLIP framework [82, 50] using the DataComp-1B dataset [30], which is one of the largest publicly available datasets with high quality. Specifically, we employ two training protocols: short schedule for fast benchmarking vision models across model and data scales, and long schedule for training best performing vision models. With the short schedule, we re-benchmark state-of-the-art vision models found on ImageNet settings for VLMs. Particularly, we select ViT [23], ConvNeXt [72], CoAtNet [18], as representatives for pure Transformer, pure ConvNet, and hybrid architecture, respectively. We combine various model scales and data scales to provide a comprehensive evaluation towards different architectures, revealing several critical findings. First, increasing data scales improves all vision models across all model sizes, while ViT scale slightly better than others in terms of model parameters. Second, the final resolution of the extracted features affects prediction performance. Third, CoAtNet performs better than ViT and ConvNeXt in general, though it is hard to scale up CoAtNet-4 to billions of data due to computational constraints.
Those findings motivate us to develop a new vision model, named ViTamin tailored for VLM. ViTamin is a 3-stage hybrid architectures, combining two stages of MBConv blocks with a final stage of Transformer blocks. This hybrid design leverages its Transformer stage to enhance data and model scalability, along with output stride of 16 to enjoy high feature resolution. As a result, ViTamin-L outshines its ViT-L/14 counterpart [30] by +2.0% zero-shot imageNet accuracy in identical OpenCLIP training scheme and identical 256 token length. When increasing feature resolution to 576 patches, ViTamin-L further attains 81.8% zero-shot imageNet accuracy, surpassing the prior art ViT-L/14 CLIPA-v2 [60] by +1.5%. In average performance across 38 datasets, it not only exceeds ViT-L/14 counterpart [60] by +1.5%, but also outperforms the larger ViT-H/14 model [60] by +0.4% while having only half parameters. When further scaling up the model size, our ViTamin-XL with only 436M parameters attains 82.9% ImageNet zero-shot accuracy, surpassing 82.0% achieved by EVA-E (i.e., EVA-02-CLIP-E/14 [94]) that has ten times more parameters (4.4B). Furthermore, we introduce an effective training scheme Locked-Text Tuning (LTT), which guides the training of vision backbone with a frozen pretrained text encoder. It enhances the small variant by +4.0% and the base variant by +4.9% without any extra cost.
Our another intriguing observation is the prevailing emphasis on data filtering over vision architecture design in VLM. For instance, while the best DataComp challenge solution [119] achieved only a +2.3% gain, our ViTamin with LTT largely improves performance by +23.3% on the same dataset size, without intensive data filtering. Finally, we introduce a suite of downstream tasks, including open-vocabulary detection and segmentation, and LMMs, for evaluating VLM-specific vision models. ViTamin outperforms the ViT-L model, enhancing detector by +3.1% on OV-LVIS and segmentor by +2.6% on average 8 datasets, and excelling across 12 LMM benchmarks. Notably, ViTamin sets a new state-of-the-art on 7 benchmarks for open-vocabulary segmentation.
We aim for our findings to encourage a reevaluation of the current limitations in VLM designs and hope that our extensive benchmarking and evaluations will drive the development of more advanced vision models for VLMs.
2 Related Work
Vision Backbone: On the ImageNet benchmark [86], ConvNets [53, 91, 96, 37, 46, 113, 88, 97, 98, 72, 128] have been the dominant networks choice since the advent of AlexNet [53]. Recently, the vision community has witnessed impressive emergence of the Transformer architecture [104], a trend that began with the widespread adoption of the ViT [23] and its subsequent developments [83, 106, 99, 26, 71, 108, 136, 62, 126, 58]. Among these, hybrid architectures [34, 93, 24, 41, 109, 18, 112, 76, 14, 103, 63, 116] combine Transformer self-attention with convolution, where CoAtNet [18] particularly obtains impressive results on ImageNet. Notably, MaX-DeepLab [107], emerged as early as 2020, successfully developed a hybrid network backbone for dense pixel predictions, where the first two stages utilize residual bottleneck blocks [37], followed by two subsequent stages employing axial attention [106]. More recently, by leveraging the design practices of a Vision Transformer, a ResNet [37] can be modernized to ConvNeXt [72], competing favorably with ViT. Along the same direction, but not limited to the ImageNet scale, our work aims to develop a novel vision model for training with billions of data [30] in the vision-language era.
Language-Image Pre-training: Language-image pre-training has seen significant advancements [50, 82, 57, 3, 120, 68, 1] with the emergence of LLMs [7, 81, 101]. The huge progress can be attributed to the pre-training on an immense scale of noisy web-collected image-text data [89, 30], much larger than the ImageNet. Notably, CLIP [82, 50] generates strong image features and excels in zero-shot transfer learning [85, 77, 31, 54, 124, 68, 137], which make it an essential role in large multi-modal model [57, 68, 13, 67]. CLIP has been improved by advanced training strategies including self-supervised learning [79, 65], efficient tuning [132, 69] and training [61, 59, 94, 110, 133]. These studies predominantly employ ViT [23] as the only vision model. As a result, the architectural design for the CLIP vision model has not been thoroughly investigated. Thus, we attempt to bridge the gap by developing a novel vision model for VLMs.
3 Method
In the section, we revisit the problem definition of CLIP and propose two training protocols (short and long schedules) on DataComp-1B (Sec. 3.1). With short schedule, we re-benchmark modern vision models found on ImageNet under the CLIP setting (Sec. 3.2). We then introduce the proposed ViTamin architecture design, motivated by the discoveries in the re-benchmarking results (Sec. 3.3).
3.1 CLIP and Training Protocols
CLIP Framework: Given a batch of $N$ image-text pairs $\{\left(I_{1},T_{1}\right),...,\left(I_{N},T_{N}\right)\}$ (where $I_{i}$ and $T_{i}$ denote image and text for $i_{th}$ pair), the objective of CLIP [82] learns to align the image embeddings $\mathbf{x}_{i}$ and text embeddings $\mathbf{y}_{i}$ for each pair. Formally, the loss function is defined as follows:
$-\frac{1}{2N}\sum_{i=1}^{N}(\underbrace{\log\frac{e^{\mathbf{x}_{i}^{T}\mathbf{y}_{i}/\tau}}{\sum_{j=1}^{N}e^{\mathbf{x}_{i}^{T}\mathbf{y}_{j}/\tau}}}_{\text{image to text }}+\underbrace{\log\frac{e^{\mathbf{y}_{i}^{T}\mathbf{x}_{i}/\tau}}{\sum_{j=1}^{N}e^{\mathbf{y}_{i}^{T}\mathbf{x}_{j}/\tau}}}_{\text{text to image }}),\vspace{-2mm}$ (1)
where $\mathbf{x}_{i}=\frac{f\left(I_{i}\right)}{\left\|f\left(I_{i}\right)\right\|_{2}}$, $\mathbf{y}_{i}=\frac{g\left(T_{i}\right)}{\left\|g\left(T_{i}\right)\right\|_{2}}$, and $\tau$ is a temperature variable. A vision model $f\left(.\right)$ and a text model $g\left(.\right)$ are trained to minimize the loss function. We focus on vision model design and use the text models from OpenCLIP [48].
Training Protocols: We employ two training protocols: short schedule and long schedule. The short schedule is designed for efficiently benchmarking vision models up to 1 training epoch on DataComp-1B [30] (i.e., 1.28B seen samples). As detailed in Tab. 2, given a descent amount of resources (e.g., 32 A100 GPUs), it takes less than two days to train a small ($\sim$25M parameters) model variant. The long schedule is designed for training the best performing models with up to 40B seen samples.
3.2 Benchmarking Vision Models in CLIP Setting
Figure 2: Benchmarking vision models under CLIP setting on DataComp-1B, including ViT (a pure Transformer), ConvNeXt (a pure ConvNet), and CoAtNet (a hybrid model). We examine their scalability in terms of both data sizes (1st row) and model scales (2nd row), and further analyze the results from the aspects of feature resolution (3rd row) and hybrid architecture (4th row).
The short schedule allows us to efficiently re-benchmark state-of-the-art vision models found on ImageNet under the CLIP setting using DataComp-1B. The experimented models are ViT [23] (a pure Transformer), ConvNeXt [72] (a pure ConvNet), and CoAtNet [18] (a hybrid model). We examine their scalability in terms of both model scales and data sizes. Each vision model has sizes varying from small ($\sim$25M parameters), base ($\sim$85M) to large ($\sim$300M), while the data sizes range from 128M, 512M to 1.28B training seen samples (1 epoch is equal to 1.28B seen samples). The metric is zero-shot accuracy on ImageNet, supplemented by the results on the 38 tasks following DataComp [30]. As shown in Fig. 2, we analyze the benchmarked results from four aspects, including data scalability, model scalability, feature resolution, and hybrid architecture. For simplicity, we use “X@Y” to denote the vision model X trained with Y seen samples. See appendix for numerical results.
Data Scalability: When training seen samples increase from 128M to 1.28B, we observe a consistent trend of improvements across all model sizes and all vision models (a1-a5). Interestingly, ViT-S/16@512M (22M parameter) attains the zero-shot performance of 53.8% on ImageNet, which is better than 45.8% by ViT-B/16@128M (86M parameter). It shows the effectiveness of training large scale data that quadrupling training seen samples can be more impactful than quadrupling the number of model parameters. Additionally, ViT-B/16@512M & @1.28B significantly boost ViT-B/16@128M from 45.8% to 60.0% (+14.2%) and 65.6% (+19.8).
$\rightarrow$ As the training seen samples increase, the performances consistently improves in all cases.
Model Scalability: When the model sizes increase, the performances of all vision models are also boosted (b1-b3). However, we observe a different gain among them (b4). For example, ConvNeXt-XL@128M brings only +1.4% gain over ConvNeXt-B@128M, while ViT-L/16@128M enhances ViT-B/16@128M by +3.6%. Given plenty of data, ViT still shows a better model scalability, especially scaling from base to large (e.g., +6.4% for ViT vs. +3.6% for both CoAtNet and ConvNeXt at 512M samples; +6.3% for ViT vs. +2.8% for CoAtNet and + 4.9% for ConvNeXt at 1.28B samples). As a result, ViT shows the best scalability.
$\rightarrow$ ViT demonstrates the most effective scaling in terms of model parameters.
Feature Resolution: Across all model scales and data sizes, ConvNeXt performs better than ViT/32 but loses its advantage to ViT/16 (c1 & c2). This trend deviates significantly from what is observed in ImageNet era, where ConvNeXt consistently outperforms ViT/16 (also see our ImageNet-scale VLM experiments in appendix). We hypothesize that, comparing to ImageNet’s object class label, the text in CLIP captures broader area of information, and thus is beneficial from higher feature resolution. Besides, ViT also benefits from using smaller patch sizes (thus high feature resolution) over larger path sizes (c3 & c4).
$\rightarrow$ The final resolution of extracted features affects the prediction performance. ViT with smaller patch sizes outperforms ViT with larger patch sizes and ConvNeXt.
Hybrid Architecture: We observe that ConvNeXt consistently lags behind ViT-{S,B}/16 and particularly ViT-L/14, suggesting that a pure ConvNet has limited capacity under the CLIP setting when presented with abundant of data (d1-d3). By contrast, CoAtNet significantly surpasses both ViT and ConvNeXt (e.g., CoAtNet-2@1.28B has a remarkable +2.9% and +3.2% gain over ViT-B/16@1.28B and ConvNeXt-B@1.28B, respectively), indicating the effectiveness of hybrid models. However, CoAtNet requires the most GPU memory; we can only train CoAtNet-4 with batch size 8k on 64 A100 GPUs, while all the other large models are trained with batch size 16k on 32 A100 GPUs. This affects CoAtNet’s scalability in large variant.
$\rightarrow$ CoAtNet surpasses ViT and ConvNeXt in general, yet it is hard to scale up CoAtNet-4 to billions of data.
3.3 Novel Vision Transformer for Vision-Language
Herein, we distill from the aforementioned observations, culminating in the proposed vision model, ViTamin (Vision TrAnsforMer for vIsion-laNguage), which notably takes the lead in the benchmarking results across all settings in Fig. 2. To introduce ViTamin, we commence by its macro-level network design (Sec. 3.3.1), followed by the micro-level block design (Sec. 3.3.2). Finally, we develop a vision model family with a simple scaling rule (Sec. 3.3.3).
Figure 3: Overview of ViTamin architecture. (a) ViTamin begins with a convolutional stem, followed by Mobile Convolution Blocks (MBConv) in stage 1 and 2, and Transformer Blocks (TFB) in stage 3. The 2D input to the stage 3 is flatten to 1D. For the macro-level designs, the three-stage layout generates the final feature map with output stride 16, similar to ViT/16 [23]. We set channels sizes for the three stages to be ($C$, $2C$, $6C$). For the micro-level designs, the employed MBConv-LN modifies MBConv [88] by using a single LayerNorm [4]. TFB-GeGLU upgrades TFB’s FFNs [104] (Feed-Forward Networks) with GELU Gated Linear Units [90]. (b) In the CLIP framework, given $N$ image-text pairs, the vision model’s output $I_{i}$ is learned to align with its corresponding text Transformer’s output $T_{i}$. Our text Transformers are the same as OpenCLIP [48]. +: Addition. *: Multiplication.
3.3.1 Macro-level Network Design
Overview: The macro-level network design of ViTamin is inspired by the ViT and CoAtNet. Specifically, on top of a simple convolutional stem (i.e., two $3\times 3$ convolutions) [18], we adopt a 3-stage network architecture, where the first two stages employ the Mobile Convolution Blocks (MBConv) [88, 43] and the third stage uses the Transformer Blocks (TFB) [104, 23]. Fig. 3 shows the overview of ViTamin. We detail the design principles below, based on the discoveries from the re-benchmarking results
Data and Model Scalability: ViT demonstrates the best scalability in terms of both model scales and data sizes. We thus opt for using Transformer Blocks in our last stage, and we stack most blocks here across different model sizes.
Feature Resolution: We tailor the network to generate high resolution feature maps in the end. Our 3-stage network design thus yields a feature map with output stride 16 (i.e., a downsampling factor of 16).
Hybrid Architecture: Similar to CoAtNet, we employ MBConv in the first two stages, resulting in a hybrid model. However, unlike CoAtNet that is constrained by its large memory usage, we propose a light-weight design of stage 1 and 2, which contain only two and four MBConv blocks.
Given the macro-level network design, we then move on to further improve the micro-level block design below.
3.3.2 Micro-level Block Design
Overview: The proposed ViTamin depends on two types of blocks: Mobile Convolution Blocks (MBConv) and Transformer Blocks (TFB). We refine each block in our model.
MBConv-LN: The Mobile Convolution Block (MBConv) [88] employs the “inverted bottleneck” design [37], starting with a first $1\times 1$ convolution to expand the channel size, followed by a $3\times 3$ depthwise convolution [44] for spatial interaction, and ending with another $1\times 1$ convolution to revert to the original channel size. Modern MBConv, as in MobileNetv3 [43], adds numerous batch normalization (BN) [49] layers and squeeze-and-excitation (SE) [45]. We adopt a simple modification by removing all BN layers and SE, and just using a single layer normalization (LN) [4] as the first layer in our block, akin to the pre-norm layer in the Transformer block, resulting in the proposed MBConv-LN. Ablation (in appendix) shows that MBConv-LN enjoys a simple design while attaining a similar performance to the original MBConv-BN-SE in MobileNetv3.
TFB-GeGLU: The Transformer Block (TFB) [104] contains two residual blocks: one with self-attention and the other with feed-forward network (FFN). We empirically discover that substituting the first linear layer with GeGLU [90], an enhanced version of the Gated Linear Unit [19] that has a 2×2$\times$2 × expansion rate, can enhance accuracy in the FFN. We denote the Transformer Block with the updated FFN as TFB-GeGLU. Ablation (in appendix) shows that TFB-GeGLU requires 12% fewer parameters than TFB due to half expansion ratio, allowing us to stack additional Transformer blocks towards deeper architectures [96, 136, 100].
3.3.3 Meta Architecture and Scaling Rule
Meta Architecture: After introducing our macro-level network and micro-level block designs, we now put everything together to form the meta architecture of ViTamin. Specifically, ViTamin is a hybrid architecture that contains only three stages, built on top of a simple convolutional stem (i.e., two $3\times 3$ convolutions). The first two stages are composed of MBConv-LN, where we stack two and four of them for stage 1 and 2, respectively. The third stage are obtained by stacking $N_{B}$ TFB-GeGLU blocks. With the meta architecture in mind, we are ready to discuss the scaling rule to generate a family of ViTamin with different model sizes.
Scaling Rule: Our scaling rule is extremely simple and straightforward, controlled by two hyper-parameters: width (i.e., the channel sizes of those three stages) and depth (i.e., $N_{B}$, the number of TFB-GeGLU blocks in stage 3). Note that our convolutional stem has the same channel size as the first stage. We define four model sizes: Small, Base, Large, and X-Large (S, B, and L variants have a similar amount of model parameters to ViT [23, 131]). We use the same channel size as ViT in our 3rd stage for each model variant. Specifically, we set the channel sizes of our three stages as $(C,2C,6C)$, where $6C=\{384,768,1024,1152\}$ for Small, Base, Large and X-Large model variant, respectivelyWe calculate the channel size for stage 1 as $1/6C$, rounding to the nearest value that is divisible by 32.. Subsequently, given the target model parameter, the value of $N_{B}$ (i.e., the number of TFB-GeGLU blocks in stage 3) can be easily found. We show the family of ViTamin-{S, B, L, XL} in Tab. 1.
ViTamin-S ViTamin-B ViTamin-L ViTamin-XL
block stride B C B C B C B C
conv-stem 2 2 64 2 128 2 160 2 192
MBConv-LN 4 2 64 2 128 2 160 2 192
MBConv-LN 8 4 128 4 256 4 320 4 384
TFB-GeGLU 16 14 384 14 768 31 1024 32 1152
Table 1: ViTamin model variants. ViTamin variants differ in the number of blocks B and number of channels C in each stage.
short schedule for benchmarking long schedule
ViTamin-S ViTamin-B ViTamin-L ViTamin-L ViTamin-XL
batch size 8k 8k 16k 90k 90k 90k
image size 224 224 224 224 256 256
# A100 GPUs 32 32 32 184 312 312
# epochs 1 1 1 10 10 30
seen samples (B) 1.28 1.28 1.28 12.8 12.8 40.0
training days 1.8 3.3 5.6 11 15 46
Table 2: Short and long training schedules on DataComp-1B.
Locked-Text Tuning for CLIP: Besides model design, we propose Locked-Text Tuning (LTT) to exploits a pretrained frozen text encoder. In light of the aligned image and text embeddings in CLIP, we leverage the pretrained text encoder from a large VLM to guide the training of image encoders of smaller VLMs. Specifically, when training other ViTamin variants (e.g., ViTamin-S and ViTamin-B), we initialize their text encoder with the one from a pretrained ViTamin-L. The text encoder is then frozen, used as a teacher to guide the training of the randomly initialized image encoder. This scheme can be considered as a way to distill the knowledge [42] from a pretrained frozen text encoder to a randomly initialized image encoder.
image num text encoder seen training training trainable params MACs ImageNet avg. 38 ImageNet VTAB retrieval
image encoder size patches depth/width samples (B) scheme dataset Image+Text (M) Image+Text (G) Acc. datasets dist. shift.
ViT-L/14 [30] 224 256 12 / 768 12.8 OpenCLIP DataComp-1B 304.0 + 123.7 77.8 + 6.6 79.2 66.3 67.9 65.2 60.8
ViT-L/14 [60] 224 256 12 / 768 12.8 + 0.5 CLIPA-v2 DataComp-1B 304.0 + 110.3 77.8 + 2.7 79.7 65.4 68.6 62.9 60.6
ViT-L/14 [60] 336 576 12 / 768 12.8 + 0.5 + 0.1 CLIPA-v2 DataComp-1B 304.3 + 110.3 174.7 + 2.7 80.3 65.7 70.2 62.5 61.1
ViTamin-L 224 196 12 / 768 12.8 OpenCLIP DataComp-1B 333.3 + 123.7 72.6 + 6.6 80.8 66.7 69.8 65.3 60.3
ViTamin-L 256 256 12 / 768 12.8 + 0.2 OpenCLIP DataComp-1B 333.4 + 123.7 94.8 + 6.6 81.2 67.0 71.1 65.3 61.2
ViTamin-L 336 441 12 / 768 12.8 + 0.2 OpenCLIP DataComp-1B 333.6 + 123.7 163.4 + 6.6 81.6 67.0 72.1 64.4 61.6
ViTamin-L 384 576 12 / 768 12.8 + 0.2 OpenCLIP DataComp-1B 333.7 + 123.7 213.4 + 6.6 81.8 67.2 72.4 64.7 61.8
ViTamin-L2 224 196 24 / 1024 12.8 OpenCLIP DataComp-1B 333.6 + 354.0 72.6 + 23.3 80.9 66.4 70.6 63.4 61.5
ViTamin-L2 256 256 24 / 1024 12.8 + 0.5 OpenCLIP DataComp-1B 333.6 + 354.0 94.8 + 23.3 81.5 67.4 71.9 64.1 63.1
ViTamin-L2 336 441 24 / 1024 12.8 + 0.5 OpenCLIP DataComp-1B 333.8 + 354.0 163.4 + 23.3 81.8 67.8 73.0 64.5 63.6
ViTamin-L2 384 576 24 / 1024 12.8 + 0.5 OpenCLIP DataComp-1B 334.0 + 354.0 213.4 + 23.3 82.1 68.1 73.4 64.8 63.7
ViTamin-XL 256 256 27 / 1152 12.8 + 0.5 OpenCLIP DataComp-1B 436.1 + 488.7 125.3 + 33.1 82.1 67.6 72.3 65.4 62.7
ViTamin-XL 384 576 27 / 1152 12.8 + 0.5 OpenCLIP DataComp-1B 436.1 + 488.7 281.9 + 33.1 82.6 68.1 73.6 65.6 63.8
ViTamin-XL 256 256 27 / 1152 40.0 OpenCLIP DataComp-1B 436.1 + 488.7 125.3 + 33.1 82.3 67.5 72.8 64.0 62.1
ViTamin-XL 336 441 27 / 1152 40.0 + 1.0 OpenCLIP DataComp-1B 436.1 + 488.7 215.9 + 33.1 82.7 68.0 73.9 64.1 62.6
ViTamin-XL 384 576 27 / 1152 40.0 + 1.0 OpenCLIP DataComp-1B 436.1 + 488.7 281.9 + 33.1 82.9 68.1 74.1 64.0 62.5
ViT-L/14 [94] 224 256 12 / 768 4.0 EVA-CLIP Merged-2B 333.3 + 123.7 72.6 + 6.6 79.8 64.9 68.9 62.8 63.3
ViT-L/14 [94] 336 576 12 / 768 6.0 EVA-CLIP Merged-2B 333.3 + 123.7 72.6 + 6.6 80.4 65.8 70.9 63.2 63.5
ViT-L/16 [133] 256 256 24 / 1024 40.0 SigLIP WebLI 316.0 + 336.2 78.1 + 19.3 80.5 65.6 70.2 62.5 61.1
ViT-L/16 [133] 384 576 24 / 1024 40.0 + 5.0 SigLIP WebLI 316.3 + 336.2 175.8 + 19.3 82.1 66.8 70.9 63.1 68.7
ViT-G/14 [48] 224 256 32 / 1280 39.0 OpenCLIP LAION-2B 1844.9 + 694.7 473.4 + 48.5 80.1 66.7 69.1 64.6 63.5
ViT-H/14 [60] 336 576 24 / 1024 12.8 + 0.5 + 0.1 CLIPA-v2 DataComp-1B 632.5 + 354.0 363.7 + 9.7 81.8 66.8 72.4 63.7 62.6
ViT-E/14 [94] 224 256 24 / 1024 4.0 EVA-CLIP LAION-2B 4350.6 + 354.0 1117.3 + 23.3 82.0 66.9 72.0 63.6 62.8
ViT-G/14 [60] 336 576 32 / 1280 12.8 + 0.5 + 0.1 CLIPA-v2 DataComp-1B 1845.4 + 672.3 1062.9 + 20.2 83.1 68.4 74.0 64.5 63.1
SoViT-400M/14 [2] 224 256 27 / 1152 40.0 SigLIP WebLI 428.2 + 449.7 106.2 + 6.6 82.0 68.1 69.5 64.8 66.8
SoViT-400M/14 [2] 384 729 27 / 1152 40.0 + 5.0 SigLIP WebLI 428.2 + 449.7 302.3 + 26.3 83.1 69.2 72.4 64.6 69.8
ViT-H/14 [27] 224 256 24 / 1024 39.0 OpenCLIP DFN-5B 632.1 + 354.0 162.0 + 23.3 83.4 69.6 69.9 67.5 68.3
ViT-H/14 [27] 378 729 24 / 1024 39.0 + 5.0 OpenCLIP DFN-5B 632.7 + 354.0 460.1 + 23.3 84.4 70.8 72.8 68.5 69.5
Table 3: Comparison with state-of-the-art models. Our models are only trained on the publicly available DataComp-1B [30]. CLIPA-v2 [60] uses an advanced progressive training scheme (from smaller images to larger ones) than the original OpenCLIP [48, 30] scheme that we follow. Other methods that use different settings are marked in gray for reference. Specifically, EVA-CLIP [94] uses EVA weights [28], better training scheme FLIP [65], and different training datasets [89, 28]. SigLIP [133] employs better sigmoid loss, stronger text encoders, and an extremely long schedule on the proprietary WebLI dataset [12] (40B for training and another 5B seen samples for fine-tuning). $\dagger$: ViT-L/14 benefits from more image tokens by using a smaller output stride 14 than 16 that we use. To have the same image tokens, we slightly enlarge the image size (e.g., $224/14=256/16$ and $336/14=384/16$). We note that all compared results are from the OpenCLIP-results that are evaluated under the same setting to ensure a fair comparison.
4 Experimental Results
In this section, we detail the implementation in Sec. 4.1, compare with the state-of-the-arts in Sec. 4.2, and deploy ViTamin to downstream tasks, including open-vocabulary detection/segmentation, and large multi-modal models in Sec. 4.3. See appendix for ablation studies.
4.1 Implementation Details
Training Strategy: We train the VLMs using OpenCLIP [48] on the public dataset DataComp-1B [30]. Tab. 2 summarizes the settings for our training schedules and model variants. We use the short schedule to benchmark vision models and conduct our ablation studies, and long schedule to train our best ViTamin-L. We closely follow the training hyper-parameter settings in OpenCLIP [48, 30]. The training and fine-tuning details are in the appendix.
Evaluation Strategy: We follow DataComp [30] to zero-shot evaluate VLMs with a testbed of 38 tasks, including ImageNet [86], 6 distribution shift tasks [105, 5, 84, 40, 39], VTAB tasks [130], WILDS tasks [52, 87], and 3 retrieval tasks [118, 11, 6].
Other Downstream Tasks: We evaluate the trained VLM in downstream tasks. For open-vocabulary detection, we exploit the F-ViT framework [111], while for open-vocabulary segmentation, we adopt the FC-CLIP framework [124] and zero-shot evaluate on multiple segmentation datasets. Finally, we evaluate VLMs in LLaVA-1.5 [67] for LMMs across multiple benchmarks. In all the cases, F-ViT, FC-CLIP, and LLaVA employ the frozen VLM backbone to effectively ablate different pretrained VLMs.
4.2 Main Results
Comparison with other State-of-the-arts: Tab. 3 summarizes the comparison between ViTamin-L and other state-of-the-art models, which exclusively employ the ViT backbone [23] but use different training schemes and datasets. For a fair comparison, we focus on the methods that use the same training data DataComp-1B [30], but still list other methods in the table for reference. For simplicity, we use “X@Z” to denote the vision model X trained with input size Z Notation @ here is slightly abused to denote the training seen samples.. ImageNet zero-shot accuracy is our main metric; other results are still reported in the table. As shown in the table, ViTamin-L@224 outperforms ViT-L/14@224 OpenCLIP [48] by +1.6%. However, ViT-L/14 benefits from more image tokens by using a smaller output stride 14 than 16 that we use (as benchmarked in the appendix). To have the same image tokens, we slightly enlarge the image size. As a result, our ViTamin-L@256 surpasses ViT-L/14@224 OpenCLIP [48] and CLIPA-v2 [60] by 2.0% and 1.5%, respectively. After fine-tuning on larger input sizes, ViTamin-L@384 and ViTamin-L@336 still performs better than ViT-L/14@336 CLIPA-v2 [60] by +1.5% and +1.3%, respectively. Impressively, with only half the parameters, our ViTamin-L attains an average of 67.2% performance across 38 datasets, exceeding the larger ViT-H/14 CLIPA-v2 model’s performance by +0.4%. Scaling up the text encoder to match the model size of the image encoder (specifically, ViTamin-L2) notably increases zero-shot ImageNet accuracy to 82.1% and average 38 datasets performance to 68.1%. Further scaling up the model parameters (i.e., ViTamin-XL) and 40 billion seen samples reaches 82.9% zero-shot ImageNet accuracy.
Locked-Text Tuning: Fig. 4 shows that our LTT improves our ViTamin-S/-B by a large margin, especially when data sizes are small. Notably, LTT lifts ViTamin-B to the next scale of model performance, surpassing ViT-L/16 by +14% in 128M samples and +1.1% in 512M seen samples. Interestingly, LTT can save 10% training budget for ViTamin-B as the text tower is completely frozen.
Figure 4: Locked-text tuning (LTT). LTT exploits a pretrained frozen text encoder, and effectively boosts the model performance.
Data Quality vs. Model Capacity: The DataComp challenge [30] underscores the role of data filtering for VLM, however, using a fixed ViT model. As shown in Tab. 4, the leading solution [119] of DataComp challenge in ICCV 2023 employed a complicated 24 filtering rules to improve the dataset quality, resulting in +2.3% gain. Surprisingly, our ViTamin-B improves the performance by a healthy margin of +12.8% accuracy, and locked-text tuning can lift the gain to +23.3%. The result highlights the importance to co-design the vision-language dataset and model.
dataset seen IN acc. avg. 38
image encoder data filtering size samp. (%) datasets (%)
leaderboard
ViT-B/32 DataComp [30] 14M 128M 29.7 32.8
ViT-B/32 SIEVE [75] 24M 128M 30.3 (+0.6) 35.3 (+2.5)
ViT-B/32 Top-1 Solution [119] 23M 128M 32.0 (+2.3) 37.1 (+4.3)
our experiments
ViT-B/32 DataComp [30] 14M 128M 29.4 31.5
ViT-B/16 DataComp [30] 14M 128M 35.8 (+6.4) 34.6 (+3.1)
ViTamin-B DataComp [30] 14M 128M 42.2 (+12.8) 38.3 (+6.8)
ViT-B/16-LTT DataComp [30] 14M 128M 43.6 (+14.2) 41.1 (+9.6)
ViTamin-B-LTT DataComp [30] 14M 128M 52.7 (+23.3) 47.2 (+15.7)
Table 4: Data quality vs. model capacity. The leaderboard results are from ICCV 2023 DataComp challenge medium filtering track.
4.3 New Suite of Downstream Tasks
The evaluations so far are mostly on classification/retrieval-based task, highlighting a lack of downstream tasks similar to those employed in the ImageNet era. Yet, in contrast to ImageNet-based vision models where downstream tasks mainly involve transfer learning for conventional detection and segmentation, VLMs excel with zero-shot capability and provides feature embeddings that are well-aligned across the vision-language domain. In light of this, we introduce a novel suite of downstream tasks aimed at the holistic evaluation of VLMs, including open-vocabulary detection and segmentation and multi-modal LLM.
pretraining OV-COCO [129] OV-LVIS [35]
image encoder dataset scheme ($\text{AP}^{\text{novel}}_{\text{50}})$ ($\text{AP}_{\text{r}}$)
ViT-L/14 DataComp-1B CLIPA-v2 36.1 32.5
ConvNeXt-L LAION-2B OpenCLIP 36.4 29.1
ViTamin-L DataComp-1B OpenCLIP 37.5 35.6
Table 5: Open-vocabulary detection. Different image encoders (ViT-L/14 by [60], ConvNeXt-L by [48]) are using the F-ViT framework [111] in a sliding window manner [125], trained on OV-COCO [129] and OV-LVIS [35]. ConvNeXt-L is marked in gray due to different pretrained dataset.
panoptic dataset (PQ) semantic dataset (mIoU)
image pretraining ADE Cityscapes MV A-150 A-847 PC-459 PC-59 PAS-21
encoder dataset scheme [134] [17] [80] [134] [134] [78] [78] [25]
ViT-L/14 DataComp-1B CLIPA-v2 24.6 40.7 16.5 31.8 14.3 18.3 55.1 81.5
ConvNeXt-L LAION-2B OpenCLIP 26.8 44.0 18.3 34.1 14.8 18.2 58.4 81.8
ViTamin-L DataComp-1B OpenCLIP 27.3 44.0 18.2 35.6 16.1 20.4 58.4 83.4
Table 6: Open-vocabulary segmentation. Different image encoders (ViT-L/14 by [60], ConvNeXt-L by [48]) are using the FC-CLIP framework [124] in a sliding window manner [125], trained on COCO [66] and zero-shot evaluated on the other datasets. ConvNeXt-L is marked in gray due to different pretrained dataset.
Open-Vocabulary Detection and Segmentation: To examine how well the trained VLMs can adapt to downstream tasks, we consider two simple yet effective frameworks F-ViT [111] and FC-CLIP [124] which utilize a frozen CLIP backbone for open-vocabulary detection and segmentation, respectively. Specifically, we consider different VLMs as plug-in frozen backbones to these frameworks, while for ViT and ViTamin that may not easily generalize to high resolution input, we extract the feature in a sliding window manner [125], with window size equal to the pre-train image size, resulting in Sliding F-ViT and Sliding FC-CLIP, respectively. Tab. 5 illustrates that ViTamin-L serves as a stronger image encoder for open-vocabulary detection, surpassing its ViT-L/14 counterpart by 1.4% and 3.1% on OV-COCO and OV-LVIS. Tab. 6 shows that ViTamin-L outperforms ViT-L/14 by 2.6% on average 3 panoptic datasets and by 2.6% on average 5 semantic datasets. Notably, surpassing prior art, ViTamin-L sets a new state-of-the-art performance across seven benchmarks for open-vocabulary panoptic segmentation and semantic segmentation.
Large Multi-modal Models: Another key application of VLMs lies in their role as vision encoders within LMMs  [57, 137, 68], as image features in VLMs that is well-aligned with text, thereby bridging the visual comprehension gap for LLMs. Specifically, we consider LLaVA-1.5 [67] as the evaluated framework. We follow [67] for all experimental settings, where the image is processed through a frozen CLIP model and a MLP projector, retaining the image as visual tokens, which are prepended to a text sequence and fed into a frozen Vicuna-v1.5-7B [16]. We run evaluation on 12 LMM benchmarks following [67], with results in Tab. 7. It should be noted that while OpenAI-trained ViT-L/14 underperforms CLIPAv2-trained counterpart by -3.7% ImageNet accuracy, it excels remarkably in LLaVA (+4.4% on VQAv2 and +4.3% on VizWiz). This highlights the need for incorporating a variety of downstream tasks to ensure a comprehensive evaluation. Surprisingly, simply replacing LLaVA’s image encoder to ViTamin-L can achieve new state-of-the-art across various benchmarks.
image training
VQAv2
GQA
VizWiz
SQA
T-VQA
POPE
MME
MMBench
MMBCN
SEED
LLaVAW
MM-Vet
encoder scheme [33] [47] [36] [74] [92] [29] [64] [70] [70] [56] [68] [127]
ViT-L/14 OpenAI 78.5 62.0 50.0 66.8 58.2 85.9 1511 64.3 58.3 58.6 65.4 31.1
ViT-L/14 CLIPA-v2 75.9 60.3 48.8 65.6 55.0 84.9 1396 60.8 54.6 54.6 60.6 28.6
ViTamin-L OpenCLIP 78.4 61.6 51.1 66.9 58.7 84.6 1421 65.4 58.4 57.7 64.5 33.6
ViTamin-L OpenCLIP 78.9 61.6 55.4 67.6 59.8 85.5 1447 64.5 58.3 57.9 66.1 33.6
Table 7: Large Multi-modal Model (LMM) performance with different VLMs. The results in 1st row originate from LLaVa-1.5 paper [67] and are marked in gray due to pretraining on OpenAI WIT dataset [82] unlike DataComp-1B [30] used by other rows. All listed models are trsained following the same settings in LLaVA-1.5 [67] with Vicuna-V1.5-7B [16], for a fair comparison. $\dagger$: image size of 384 rather than the default 336.
5 Conclusion
In this work, we build an evaluation protocols of modern vision models in VLM and re-benchmark them under CLIP setting. We examine vision models from four aspects of data scalability, model scalability, feature resolution and hybrid architecture. The four pillars motivate us to propose ViTamin, which not only competes favorably with ViT in zero-shot ImageNet accuracy and average 38 dataset accuracy, but also achieves the state-of-the-art on 22 downstream tasks covering open-vocaburary detection and segmentation and large multi-modal models. We hope that our design practices will drive the development of more advanced vision models for VLMs.
Acknowledgement: We thank Haichao Yu and Zhanpeng Zeng for the discussions about DataComp challenge and micro-level block design, respectively. This work was supported in part by ONR N00014-23-1-2641.